OpenStudy (anonymous):

1. The pet store has a treat bar that sells treats for dogs. The price for 3 lb of Carob Cookies is $8.25. The price for is 5 lb of Ginger Bears is $12.75. One customer spent $21.60 on dog treats. He purchased 6 lb of Carob Cookies. How many pounds of Ginger Bears did he purchase? Write an equation for this situation.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino Help! :3

OpenStudy (michele_laino):

here we have: cost of one lb of cookies = 8.25/3=...?

OpenStudy (anonymous):

2.75

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

cost of one lb of Ginger Bears = 12.75/5=...?

OpenStudy (michele_laino):

what is 12.75/5=...?

OpenStudy (anonymous):

2.55

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@nincompoop plz help me

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

can someone plz help me!

OpenStudy (anonymous):

@Preetha plz help me

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

plz someone help me

OpenStudy (anonymous):

@nincompoop @misty1212 @perl @Preetha

OpenStudy (anonymous):

please this is due tomorrow